After completing an inventory of three warehouses, a golf club shaft manufacturer described its stock of 12,246 shafts with the percentages given in the table. Suppose a shaft is selected at random from the 12,246 currently in stock, and the warehouse number and type of shaft are observed Type of Shaft Regular 19% 14% 7% Stiff 8% 20% 18% Extra Stiff 990 Warehouse 2 096 Find the probability that the shaft was produced in a warehouse other than warehouse 1 0.40 0.64 0.36 0.86
Answer #1

Percentage of shafts that were produced in warehouse 2 = (14 + 20 + 5)% = 39%

Percentage of shafts that were produced in warehouse 3 = (7+ 18 + 0)% = 25%

The probability that the shaft was produced in a warehouse other than warehouse 1

= 0.39 + 0.25

= 0.64 (ans)